# The German Tank Problem with Multiple Factories

### 21 Mar 2024 v1

@@ Abstract.
During the Second World War, estimates of the number of tanks deployed by Germany were critically needed. The Allies adopted two methods to estimate this information: espionage and statistical analysis. The latter approach was far more successful and is as follows: assuming that the tanks are sequentially numbered starting from 1, if we observe $k$ serial numbers from an unknown total of $N$ tanks, with the highest observed number being $M$, then the best linear unbiased estimator for $N$ is $M(1+1/k)-1$. This is now known as the German Tank Problem and serves as an outstanding demonstration of the practical applications of mathematics and statistics in the real world. We present a natural generalization of this statistical approach.
@@ 2.1. Proof GTP and GTP-UM are the Minimum-Variance Unbiased Estimators
The Lehmann-Scheffé theorem states that any estimator that is unbiased for a given unknown quantity, and that depends on the samples only through a complete, sufficient statistic, is the unique best unbiased estimator of that quantity - the MVUE (minimum-variance unbiased estimator). See [Wi] for details.
@@ 3.2. Limiting Behaviour of $P$ as $N,l,k\to\infty$
Observe that $|\rho|$ is bounded from above by $2l^{-1}$ and by the squeeze theorem we get that $\rho\to 0$ as $l\to\infty$. Next consider $k=\alpha l^{{1}/{c}}$ where $\alpha=A^{-1}$. We ignore the floor function as we can always bound the floor function $l=\lfloor Ak^{c}\rfloor$ from above and below in the limit by some continuous functions of the form $\hat{l}=Bk^{c}$. We want to show that $\sigma\to 0\text{ as }l,k\to\infty\ \ \forall A,c>0$, so that we can apply Chebyshev’s inequality [Wi] as the variance will tend to 0.
@@ 4. Restricting to Fixed, Known Gaps
For general $k$, we can approximate $\hat{N}$. The only issue arises in calculating the final sum in (4.5). For this, we can expand the binomial coefficient and then use the leading term for Faulhaber’s formula for the sum of $k^{\text{th}}$ powers of natural numbers [Wi] to approximate the sum for large $N$. For the following, take $k\ll lN$:
$\begin{split}\frac{G}{\binom{lN}{k}}\sum_{t=1}^{l}\binom{tN}{k}\ &=\ \frac{Gk!}{\binom{lN}{k}}\sum_{t=1}^{l}tN\left({tN-1}\right)\cdots\left({tN-{k-1}}\right)\\ &\approx\ \frac{G}{\left({lN}\right)^{k}}\sum_{t=1}^{l}\left({t^{k}N^{k}}\right)\\ &\approx\ \frac{GN^{k}}{\left({lN}\right)^{k}}\left({\frac{1}{k+1}l^{k+1}+\frac{1}{2}l^{k}}\right)\\ &=\ G\left({\frac{1}{k+1}l+\frac{1}{2}}\right).\\ \end{split}$ (4.9)

### 11 Apr 2024 v2

@@ Abstract.
During the Second World War, estimates of the number of tanks deployed by Germany were critically needed. The Allies adopted two methods to estimate this information: espionage and statistical analysis. The latter approach was far more successful and is as follows: assuming that the tanks are sequentially numbered starting from 1, if we observe $k$ serial numbers from an unknown total of $N$ tanks, with the highest observed number being $M$, then the best linear unbiased estimator for $N$ is $M(1+1/k)-1$. This is now known as the German Tank Problem [Wi] and serves as an outstanding demonstration of the practical applications of mathematics and statistics in the real world. We present a natural generalization of this statistical approach.
@@ 2.1. Proof GTP and GTP-UM are the Minimum-Variance Unbiased Estimators
The Lehmann-Scheffé theorem states that any estimator that is unbiased for a given unknown quantity, and that depends on the samples only through a complete, sufficient statistic, is the unique best unbiased estimator of that quantity - the MVUE (minimum-variance unbiased estimator). See [CB] for details.
@@ 3.2. Limiting Behaviour of $P$ as $N,l,k\to\infty$
Observe that $|\rho|$ is bounded from above by $2l^{-1}$ and by the squeeze theorem we get that $\rho\to 0$ as $l\to\infty$. Next consider $k=\alpha l^{{1}/{c}}$ where $\alpha=A^{-1}$. We ignore the floor function as we can always bound the floor function $l=\lfloor Ak^{c}\rfloor$ from above and below in the limit by some continuous functions of the form $\hat{l}=Bk^{c}$. We want to show that $\sigma\to 0\text{ as }l,k\to\infty\ \ \forall A,c>0$, so that we can apply Chebyshev’s inequality [Mil] as the variance will tend to 0.
@@ 4. Restricting to Fixed, Known Gaps
For general $k$, we can approximate $\hat{N}$. The only issue arises in calculating the final sum in (4.5). For this, we can expand the binomial coefficient and then use the two leading terms from Faulhaber’s formula for the sum of $k^{\text{th}}$ powers of natural numbers (see [CG] for derivation and more details) to approximate the sum for large $N$. For the following, take $k\ll lN$:
$\begin{split}\frac{G}{\binom{lN}{k}}\sum_{t=1}^{l}\binom{tN}{k}\ &=\ \frac{G}{k!\binom{lN}{k}}\sum_{t=1}^{l}tN\left({tN-1}\right)\cdots\left({tN-\left({k-1}\right)}\right)\\ &\approx\ \frac{G}{\left({lN}\right)^{k}}\sum_{t=1}^{l}\left({t^{k}N^{k}}\right)\\ &\approx\ \frac{GN^{k}}{\left({lN}\right)^{k}}\left({\frac{1}{k+1}l^{k+1}+\frac{1}{2}l^{k}}\right)\\ &=\ G\left({\frac{1}{k+1}l+\frac{1}{2}}\right).\\ \end{split}$ (4.9)